Which of the following is NOT a true statement about Price to Earnings ratios?

P/E ratios are useful for comparing companies in the same sector.

P/E compares a company's market valuation with the income it is actually generating.

Stocks with higher forecast earnings growth will usually have a lower P/E.

With trailing P/E, the earnings per share is based on the most recent 12 month period.

Which of the following is a true statement about using Return on Equity (ROE) to analyze a company?

All of these answers

If a company increases the debt percentage of its capital structure, its ROE will decrease.

If a company's net margin increases, it results in a lower overall ROE.

ROE measures how efficient the company is at generating profits from the funds invested in it
Answer #1

P/E ratios are useful for comparing companies in the same sector. - TRUE As P/E ratios varies from industry to industry P/E ratios are best for comparing companies in same sector

P/E compares a company's market valuation with the income it is actually generating.- TRUE As finding P/E ratio of a company requires Market Share Price of company by EPS of the company

Stocks with higher forecast earnings growth will usually have a lower P/E. - FALSE As companies that grow faster often tend to have higher P/E ratio as investors chase the stock price of these companies due to high growth expectations pushing P/E higher

With trailing P/E, the earnings per share is based on the most recent 12 month period. - TRUE The trailing P/E is calculated by dividing price by latest 12 month EPS

Now the Du Ponts ROE equation can be given as

ROE = Net Margin * Asset Turnover * Financial Leverage

ROE = \frac{Net Income}{Sales} * \frac{Sales}{Assets} * \frac{Assets}{Shareholder's Equity}

If a company increases the debt percentage of its capital structure, its ROE will decrease.- FALSE Increased debt increases the financial leverage in a company which during normal or boom times can results in exponential profit returns whereas during recessions can result in losses as well hence Increased debt favors ROE during boom times but hurts ROE during recessions.

If a company's net margin increases, it results in a lower overall ROE. - FALSE As can be clearly seen from equation increase in Net margin increases ROE

ROE measures how efficient the company is at generating profits from the funds invested in it - TRUE ROE calculates how the company is using the funds of its investors to generate net income